### v1

# Unexpected fault activation in underground gas storage. Part II: Definition of safe operational bandwidths

Abstract
Underground gas storage is a versatile tool for managing energy resources and addressing pressing environmental concerns. While natural gas is stored in geological formations since the beginning of the 20th century, hydrogen has recently been considered as a potential candidate toward a more flexible and sustainable energy infrastructure. Furthermore, these formations can also be used to sequester environmentally harmful gases such as CO2 securely. When such operations are implemented in faulted basins, however, safety concerns may arise due to the possible reactivation of pre-existing faults, which could result in (micro)-seismicity events. In the Netherlands, it has been recently noted that fault reactivation can occur “unexpectedly” during the life of an underground gas storage (UGS) site, even when stress conditions are not expected to cause a failure. The present two-part work aims to develop a modeling framework to investigate the physical mechanisms causing such occurrences and define a safe operational bandwidth for pore pressure variation for UGS operations in the faulted reservoirs of the Rotliegend formation, the Netherlands. In this follow-up paper, we investigate in detail the mechanisms and crucial factors that result in fault reactivation at various stages of a UGS project. The mathematical and numerical model described in Part 1 is used, also accounting for the effect of geochemical dissolution on reservoir and caprock weakening. The study investigates the risks of fault activation caused by the storage of different fluids for various purposes, such as long-term CO2 sequestration, CH4 and N2 injection and extraction cycles, and N2 permanent storage. The results show how geological configuration, geomechanical properties, and reservoir operating conditions may increase the risk of fault reactivation at various UGS stages. Furthermore, the analysis indicates that reservoir pressure near critical levels and reactivation during primary production may significantly contribute to potential fault instabilities. In light of these findings, operational guidelines for improving secure and effective storage operations are thereby presented.
keywords:
Fault reactivation , Underground Gas Storage , Hydrogen Underground Storage , Carbon Capture and Sequestration , Nitrogen Storage , Safety guidelines
journal: Geomechanics for Energy and the Environment
[…]
[UNIPD]organization=Department of Civil, Environmental and Architectural Engineering, University of Padua, city=Padua, postalcode=35131, country=Italy
[…]
[M3E]organization=M3E S.r.l., city=Padua, postalcode=35121, country=Italy
[…]
[UNIRO]organization=Department of Earth Sciences, Sapienza University of Rome, city=Rome, postalcode=00185, country=Italy
[…]
[POLITO]organization=Department of Environment, Land and Infrastructure Engineering, Polytechnic of Turin, city=Turin, postalcode=10129, country=Italy
1 Introduction
The development of underground gas storage sites has been crucial in the global energy infrastructure since the early 20th century. The first operational underground storage system (USS) site dates back to 1915 (Al-Shafi et al., 2023). Over time, this technology has expanded with more than 600 facilities worldwide (Foulger et al., 2018) to effectively manage gas supply and demand on both seasonal and daily bases. While USS has typically focused on storing natural gas (UGS), it possesses the potential to store a variety of gases for diverse purposes, thus becoming a flexible tool for managing energy resources and addressing environmental concerns. This is especially important as Europe moves toward a net-zero greenhouse gas emission energy system (European Commission and Directorate-General for Climate Action, 2019). The scientific community has shown a growing interest in underground hydrogen storage (UHS) as an alternative to natural gas. Having the potential to accommodate significant volumes, ranging from tens of millions of cubic meters in caverns to potentially billions of cubic meters in depleted gas fields (Groenenberg et al., 2020), hydrogen versatility and absence of carbon emissions during production and utilization make it a more sustainable option in comparison to CH4. Additionally, carbon capture and storage (CCS) initiatives have identified underground storage of CO2 as a crucial component in mitigating greenhouse gas emissions and contrasting climate change.
However, despite the widespread use of USS, inherent risks are associated with the potential reactivation of existing geological fractures (Ellsworth, 2013; Foulger et al., 2018; Keranen and Weingarten, 2018) and consequent induced seismicity. Although such events are statistically rare (Foulger et al., 2018), they demand significant attention due to their social and economic implications. In the Netherlands, natural gas (i.e., CH4) is currently stored in four UGS facilities: Bergermeer, Alkmaar, Norg, and Grijpskerk. They are located in the Carboniferous-Rotliegend formation in northern Europe, which is considered one of the most extensively explored petroleum systems worldwide (Gautier, 2003). Over the past few decades, significant seismic activity related to induced seismicity has been observed in this region (van Wees et al., 2014; Uta, 2017), particularly in three of the four UGS facilities (i.e., Bergermeer, Norg, and Grijpskerk). Recently, the Netherlands have also focused the attention on the feasibility of storing different gases in depleted gas fields and salt caverns as a crucial measure in the ongoing energy transition, by applying the same technology used in UGS to offshore CO2 storage through the Porthos project (https://www.porthosco2.nl), while research on hydrogen storage is in progress and primarily focused on salt caverns (UHS), although the potential for storage in abandoned gas fields is also being considered for the future. Additionally, nitrogen, which is utilized to convert high-calorific gas to low-calorific gas for residential heating and cooking purposes in Dutch households, is presently stored in a salt cavern near Heiligerlee (Ministry of Economic Affairs and Climate Policy, 2021; Muntendam-Bos et al., 2022).
Given the future role that USS will play, it is crucial to understand why faults can become active and which mechanisms are potentially prone to lead to greater instability. During the cyclic or more permanent storage of these fluids, fault reactivation can occur when the natural stress regime on a fault surface is altered due to changes in pore pressure $P$ within a reservoir. The interaction between human activities at depth and the existing stress conditions determines the initiation, amount of slip, and extent of fault reactivation (Segall et al., 1994; Hettema et al., 2000; Candela et al., 2019). From a physical point of view, a fault remains locked if the applied shear stress does not exceed the frictional resistance to sliding, and this concept is well understood in terms of the effective stress principle, expressed by the equation $\tau<\tau_{L}=c+\mu(\sigma_{n}-P_{f})$ (Jaeger et al., 2007). Indeed, various factors, such as shear stress ($\tau$, $\tau_{L}$), cohesion ($c$), friction coefficient ($\mu$), normal compressive stress ($\sigma_{n}$), and fluid pressure ($P_{f}$), influence the fault behavior and susceptibility to seismic activity. Changes in these factors can occur over different timescales.
On one hand, short-term processes can be interpreted through a geomechanical perspective involving injection/extraction activities lasting decades. In this context, most human-induced seismic events can be easily explained, as they occur due to fluid injection causing pressure to rise beyond the initial level, ultimately leading to the shear stress on the fracture surface to achieve its limit strength (Walsh and Zoback, 2016; Franceschini et al., 2024). However, a subset of recorded events cannot be accounted for by this mechanism. These events, identified as “unexpected” seismic events, occur when the pressure falls within the range typically experienced during primary production. In this scenario, fault reactivation occurs not only during primary production or gas storage at pressures exceeding the natural value $p_{i}$ (Deflandre et al., 2018; Cesca et al., 2014; Zhou et al., 2017), but also during cushion gas injection or producing and storing phases with pore pressures below the natural value but above a minimum pressure (Hager and Toksoz, 2009; Kraaijpoel et al., 2013; TNO, 2015; Nederlandse Aardolie Maatschappij BV, 2016). A typical time-behavior of the fluid pressure variations expected in UGS and UHS applications is sketched in Figure 1. The recorded seismic events are unexpected because the stress state should have already been experienced by the reservoir and surrounding faults during previous production phases. Unraveling the underlying mechanisms behind these “unexpected” seismic events is crucial for ensuring the safe and efficient operation of underground gas storage facilities.
On the other side, long-term processes, spanning hundreds or even thousands of years, primarily involve fluid-rock interactions that can alter fault rock fabric, mineralogy, frictional properties, and cohesion. Presently, the understanding of the specific impacts of different gases on the mechanical properties of faults and rock is poorly understood, but the potential risk associated with their storage has to be considered. Indeed, the feasibility of the new technologies, in terms of performance, efficiency, and safety of the storage, greatly depends on fluid-rock interaction.
In this follow-up paper of Franceschini et al. (2024), we conduct an investigation into the phenomenon of “unexpected” seismic events in underground gas storage facilities. By understanding the factors contributing to such events, we can establish enhanced guidelines and best practices for operating underground gas facilities during different storage phases. We present a comprehensive study to re-assess the fundamental geomechanical causes of seismicity and specific risk factors, including the influence of geochemical dissolution effects on reservoir and caprock weakening. This work has three specific aims. The first one is to improve the understanding of the physics-based processes behind the induced seismicity during both cycling and permanent gas storage. The second aim involves investigating which are the roles of the Rotliegend reservoirs peculiarities, i.e., the graben-fault bounds, and the compartmentalization into offset blocks by intra-field faults on favoring fault reactivation. Lastly, the third aim is to propose, based on the modeling outcomes, valuable insights into the development of general guidelines for safe operational bandwidths for reservoir storage conditions during pressure cycling (and noncycling) of underground storage of CH4, CO2, H2, and N2 in depleted gas fields. These guidelines are intended to serve as a valuable tool, contributing to the effectiveness of risk assessment and mitigation strategies in underground gas storage operations. A few preliminary outcomes regarding UGS are already reported in Teatini et al. (2019, 2020) and Franceschini et al. (2024).
The analysis that follows is accomplished by using the mathematical and numerical framework developed in Part I of this work (Franceschini et al., 2024), where a Finite Element (FE) geomechanical simulator is implemented with a Lagrange multiplier-based treatment of the frictional contact conditions for reproducing the fault behavior. The simulator is supplemented also with visco-elasto-plastic constitutive laws, e.g. Isotton et al. (2019), and fault activation is ruled by the Coulomb frictional criterion. Pressure change within the faults, variation of Coulomb’s parameters due to slip-weakening, and the rheology of the caprock are properly accounted for. Because of the main aims listed above, the modeling study is not focused on a specific UGS field, but on a conceptual reservoir and fault system that realistically represents the main geologic features of the Rotliegend formation system.
The paper is organized as follows. The typical setting of the Rotliegend reservoirs and a few examples of seismicity recorded during UGS are initially presented. Then, fluids under investigation are described, and the modeling set-up is outlined. The various scenarios are then developed to understand the geomechanical behavior of the faulted system, first analyzing the parameters involved in UGS and then comparing the scenarios using different fluids. Modeling results are presented, and the mechanisms responsible for fault reactivation during the different phases are identified. The modeling outcomes are discussed, highlighting the peculiarities of the storage scenarios with respect to induced seismicity during primary production as addressed in previous studies, ranking the natural features and anthropogenic factors more prone to cause fault reactivation, and illustrating general operational guidelines to reduce the probability of reactivation occurrences.
Figure 1: Sketch of the main stages during the reservoir lifespan in terms of pressure change over time for CH4, CO2, H2 and N2: Primary Production (PP) for CH4 production, Storage (ST) for CO2 and N2 storage, Cushion Gas Injection (CGI) for H2 and CH4 injection, and Underground H2 Storage (UHS) and Underground Gas Storage (UGS) cycles for seasonal pressure fluctuations during CH4/H2 storage (UGS-S/UHS-S) and production (UGS-P/UHS-P).
2 Induced seismicity in Rotliegend UGS reservoirs
The Rotliegend formation stretches over a large region covering north-western Germany, the Netherlands, with an off-shore portion that extends and includes part of the UK. The Rotliegend sandstones have been attributed to deposition in a variety of arid, terrestrial environments, among which ephemeral fluvial (wadi) systems, various types of eolian deposits, desert-lake environments, and adjacent sabkhas are dominant. The sealing nature of Zechstein salt is the main reason for the existence of the large amount of traps that sealed gas migrated from the underlying coals and carbonaceous shales of the Carboniferous Coal Measures for approximately 150 million years (Gautier, 2003). The Rotliegend unit spans a depth range between 2000 and 4700 m and it is generally made of permeable sandstone (average 100-200 mD), with high net-to-gross and porosity (average 15-20%). The thickness of the reservoir sands is generally over 50 m and rarely exceeds 300 m. Faulting is the peculiarity of the geologic unit, with the majority of the faults that are Triassic-Tertiary features nucleated on Mid Palaeozoic fault patterns and reworked during the Variscan orogeny. The Mesozoic rifting phase established the general extensional normal fault pattern (De Jager, 2007).
As a consequence, many fields are formed by a number of rhomboid-shaped dipping fault blocks, positioned one next to each other. The trapping mechanism is structural, by the juxtaposition of the Rotliegend reservoir blocks against the thick evaoporitic Zechstein formation. Moreover, the hydraulic connectivity between the various blocks is extremely poor, causing a clear compartmentalization of the reservoirs. Fault seal and, secondarily, stratigraphic barriers and diagenesis, are the causes of this compartmentalization (van Hulten, 2010). The leftomost frame in Figure 2 sketches as a representative examples the structural map of the Norg UGS reservoir.
Refer to caption Refer to caption
Figure 2: Left: Map of the Norg UGS site (in blue) with traces of the bounding faults (black lines) and localization of the recorded seismic events. Right: pressure evolution in time at the Grijpskerk UGS field. The location of the two seismic events in 1997 and 2015 are reported (modified from TNO report (2015)).
The reservoir is bounded by normal faults with a significant (up to a 250 m) throw and consist of a few compartments also separated by internal faults. The gas fields are 2500-3500 m deep, with the Rotliegend reservoir rock characterized by an average net thickness of 150-200 m. As usual for UGS plants, the reservoir is a former partially-depleted gas field that was converted to underground gas storage after a primary production (PP) period. PP is generally followed by a cushion gas injection (CGI) phase when fluid pressure in the reservoir is almost risen to the original value, and then by the UGS injection and production cycles. The rightmost frame of Figure  2 summarizes the evolution of the average pressure in Grijpskerk, another representative UGS field. It must be highlighted that pressure differences over reservoir blocks can be as high as about 7 MPa. Gas is injected at temperatures similar or slightly below those at the reservoir level, thus minimizing thermal stresses.
The timing and location of the seismic events recorded in the vicinity of the UGS reservoirs is also shown in Figure 2. In Norg, the time of the first earthquake coincided with low pressure level during PP. The second seismic event occurred at the end of the CGI phase, with the pressure in the reservoir close to the original pressure level. In Grijpskerk, a first seismic event took place at the end of PP in 1997, and a second one much later, in 2015, almost at the end of a UGS production phase (TNO, 2015). Four seismic events with a magnitude $M\in[3.0,3.5]$ were detected in 1994 and 2001 during PP in the Bergermeer reservoir. The events were located at the tip of the central fault separating the two main blocks (TNO, 2015). A down-hole micro-seismic array characterized by a magnitude of completeness below $M=0$ was established during the CGI phase. The array recorded a large number of micro-seismic events ($M<0$) on the faults at the reservoir depth and a main event with $M=0.9$ along the central fault when the pressure difference between the two blocks peaked about 4 MPa at the beginning of 2013 (Baisch et al., 2016).
3 Chemo-mechanical effects by CO2, H2 and N2 on rocks and faults
3.1 Effect on rock and caprock
The interactions between gas, brine, and rock can lead to significant changes in the reservoir and fault properties, which, in turn, can greatly impact the performance, efficiency, and safety of the storage process. The impact of gas injection on the mechanical response of sedimentary rocks remains uncertain to date. Existing research predominantly focuses on CO2 injection (Kim and Makhnenko, 2022), with limited investigation into N2 injection (Fuchs et al., 2019). There is some ongoing development regarding H2 injection into sedimentary rocks. The latest comprehensive review on H2 injection and withdrawal has been compiled by Corina et al. (2022). From the available literature on CO2-fluid-rock interactions, e.g. Rohmer et al. (2016) and Peter et al. (2022), two contrasting viewpoints emerge: CO2 can either have a negligible influence on the rock mechanical parameters, or degrade them because of mineral dissolution and modifications in pore size distribution. A stable condition is often observed when the mineral composition and pore structure remain unchanged during CO2 injection (Rimmelé et al., 2010; Bolourinejad and Herber, 2015), with the main elastic properties largely unaffected by the CO2 presence (Mikhaltsevitch et al., 2014; Hu et al., 2016). On the other hand, in cases where a weakening effect is observed after CO2 injection, the mechanical strength of the rock decreases, leading to increased permeability and reduced rock stiffness. The dissolution of carbonate minerals, particularly calcite, reduces the cohesion of the rock matrix and weakens grain-to-grain contacts (Kim and Makhnenko, 2022). This effect has been observed in experiments on sedimentary rocks like sandstones and carbonates. Generally, carbonate-rich rocks are more susceptible to mechanical weakening due to CO2-fluid-rock interactions.
While several experimental evidences support these findings, uncertainties still persist in the results. For instance, CO2-acidified brine injection can precipitate minerals, potentially increasing the strength and stiffness of the rock depending on the specific material and its mineral composition. CO2 can react with certain minerals like calcite, inducing mineral precipitation and cementation, thereby enhancing the mechanical integrity of the rock (Espinoza et al., 2018). Sandstones, for example, may experience both increase and decrease in the deformation moduli (Marbler et al., 2012). The variability in results regarding the effects on rock stiffness and strength highlights the need for additional experimental data to address existing uncertainties. Regarding the geochemical effects of CO2 injection in the Upper Rotliegend formation, Bolourinejad and Herber (2015) reported an increase in permeability ranging from 10% to 30%. This suggests changes in porosity and deformation moduli, with potential deformation moduli variation of $\pm$30%.
Regarding N2 and H2 injection, there is a limited investigation. N2 is primarily used for determining rock porosity or as a control gas (Fuchs et al., 2019). Therefore, the exposure of rocks to N2 is expected to have negligible effects. In the case of H2 injection, quantitative data is lacking, but qualitative rock degradation is expected. Hydrogen injection in UHS porous reservoirs may trigger the dissolution of carbonates and sulfates, as well as grain crushing and local compaction, altering the reservoir porosity and mechanical and flow properties (Al-Yaseri et al., 2023). The response will be time-dependent, with pH equilibration to higher values potentially leading to further compaction and stiffening of the reservoir. However, this effect is likely to stabilize after a few cycles. Consequently, the poroelastic response of the reservoir will heavily depend on the chemical environment in the early stages of operations and become more stable and predictable in later stages.
3.2 Effect on faults
From a goechemical point of view, gas injection may also affect faults, which represent the key factors responsible for fluid migration and containment. Gas injection can affect fault permeability, reactivation potential, and overall stability, which could lead to unintended fluid pathways or seismic activity. Concerning the CO2, during the storage phase over long periods (on the order of thousands of years), fluid-rock interactions can lead to changes in fault rock fabric and mineralogy. This may result in fault frictional properties and cohesion alterations due to dissolution and cementation processes (Collettini et al., 2008). However, laboratory friction tests comparing fault rocks exposed to CO2 for several hours to samples taken from natural analogs exposed to CO2 over millions of years, show limited friction reduction due to CO2 exposure (Samuelson and Spiers, 2012). Moreover, no specific laboratory studies show how CO2-enhanced dissolution and re-precipitation processes may impact on fault cohesion. Based on the literature, the chemical effects on fault mechanical properties after CO2 injection are considered as negligible.
Finally, there is a lack of documentation on the effect of H2 and N2 injection on faults. Experience with underground hydrogen storage in porous geological formations is minimal, with practical applications restricted to the storage of town gas, i.e., a mixture of gases containing 25-60% hydrogen, along with smaller amounts of CH4, CO, and CO2 (Heinemann et al., 2021). Injecting hydrogen into a porous reservoir can change the chemical equilibrium of the formation pore water, leading to fluid-assisted grain-scale processes such as cement dissolution, clay mineral sorption/desorption within grain boundaries, stress corrosion cracking, dissolution-precipitation, and/or intergranular frictional slip (Heinemann et al., 2021). Although these grain-scale mechanisms are well-studied, little is known about the specific effects of hydrogen and nitrogen on their rates.
4 Model set-up
We use the mathematical and numerical framework developed, discussed and tested in Part I of this work (Franceschini et al., 2024). The modeling approach consists of one-way coupled Finite Element (FE) hydro-poromechanical simulations in fractured geological media. For all the details, the reader is referred to Franceschini et al. (2024).
Refer to caption Refer to caption
Figure 3: On the left: Plain view of the conceptual model. On the right: vertical sections of the conceptual model along the trace A-A and B-B.
4.1 Conceptual reference model
The domain geometry conceptually reproduces the compartmentalization commonly observed in the UGS fields of the Rotliegend formation and already sketched by Franceschini et al. (2024). The reservoir compartments are approximately located at the center of a 30$\times$30$\times$5 km domain (Figure 3). The reservoir is 200-m thick, 2-km deep, and composed by two adjacent 2$\times$2 km blocks. The two reservoir compartments are separated by a fault (F3) and confined by two sets of orthogonal faults (F1-F2 along the y-axis, F4-F5 along the x-axis). Depending on the sealing properties of fault F3, the two compartments can have a partial hydraulic connection. The faults extend from 3-km to 1.6-km depth and terminate in the Zeichestein salt formation sealing the reservoir on top. Faults F1 and F2 have a dip angle of ±$\pm 85^{\circ}$, while faults F4 and F5 are vertical. The dip angle of fault F3 can vary from $-65^{\circ}$ to $+65^{\circ}$. Block 2 can be moved up and down, with a maximum offset of 200 m with respect to the Block 1.
The 3D domain is discretized by 8-node hexahedra, with 253,165 nodes and 236,208 elements. A finer discretization is used between 1800 and 2200 m depth. Here, the reservoir element charateristic size is 100$\times$100$\times$20 m. The fault system embedded in the 3D grid is discretized by 5,215 zero-thickness contact elements. Standard boundary conditions with zero displacements on the outer and bottom surfaces are prescribed, whereas the top surface, representing the ground, has a stress-free condition. Figure 4 shows a view of the discretized domain with the embedded fault system.
Figure 4: Axonometric view of the 3D computational grid used for the geomechanical simulation and the embedded 2D grid for the fault system.
The setting of the reference scenario is established on the basis of the hydro-geomechanical properties of the Rotliegend formation. The fault F3 is kept vertical, with no offset between the reservoir blocks. Overburden, caprock, reservoir, and underburden are assumed to behave elastically, with the properties reported in Table 1. Faults are governed by Coulomb’s failure criterion with cohesion $c=2$ MPa and static friction angle $\varphi_{s}=30^{\circ}$. A linear slip-weakening is also possible with a reduction to the dynamic friction angle $\varphi_{d}$ at a sliding equal to $d_{c}$ (Franceschini et al., 2024). As to the undisturbed stress regime, we assume the vertical effective stress, $\sigma_{v}$, to be a principal component. The other two horizontal principal components of the effective stress tensor, $\sigma_{H}$ and $\sigma_{h}$, are characterized by confinement factors $M_{1}$=$\sigma_{h}$/$\sigma_{v}$=0.74 and $M_{2}$=$\sigma_{H}$/$\sigma_{v}$=0.83, with $\sigma_{h}$ orthogonal to F1, F2, and $\sigma_{H}$ orthogonal to F4 and F5. The value of $\sigma_{v}$ is computed from the formation density, assuming a hydrostatic pore pressure distribution.
4.2 Sensitivity analysis
Layer Density [kg/m3] Young Modulus [GPa] Poisson ratio [-]
Overburden 2200 10.0 0.25
Zechstein Salt 2100 40.0 0.30
Reservoir (Upper Rotliegend) 2400 11.0 0.15
Underburden 2600 30.0 0.20
Table 1: Formation-dependent geomechanical parameters.
A sensitivity analysis has been developed within the range of some geological and geomechanical quantities to investigate the configurations that are more likely to generate “unexpected” fault reactivations during the injection and storage of different kinds of fluid. The analysis addresses the role of: (i) the geological setting, by varying geometry and properties of faults and reservoir rock, (ii) the poro-elastic stress change with respect to the natural stress regime, and (iii) the space and time pore pressure gradients in the UGS formation. The analysis has been carried out in two stages:
  1. 1.
    one parameter at a time is modified with respect to the reference scenario. This enabled a thorough analysis of the individual influence of each parameter. The considered variables and scenarios are summarized in Table 2;
  2. 2.
    simultaneous variation of more than one parameter, as identified from Stage 1. The selected combination settings were tested with different injected fluids. Chemical effects were also accounted for by either weakening or hardening the Young modulus by 30% according to the type of injected fluid (Bolourinejad and Herber, 2015).
Notice that the objective of this analysis is to reproduce realistically the actual settings of the Rotliegend formation, rather then to investigate “extreme” conditions.
Paper section Scenario Parameter/mechanism
5.2 1 reference
5.3.1 2a Fault F3 with dip angle $\delta=+65^{\circ}$
2b Fault F3 with dip angle $\delta=-65^{\circ}$
2c Block 2 moved down by the offset $o=100$ m
2d Block 2 moved down by the offset $o=200$ m
5.3.2 3a Fault cohesion $c=0.5$ MPa
3b Fault static friction angle $\varphi_{s}=20^{\circ}$
3c Linear slip-weakening with $\varphi_{d}$ = 10 and $d_{c}=2$ mm
3d Linear slip-weakening with $\varphi_{d}$ = 20 and $d_{c}=20$ mm
4a Reservoir Young’s modulus $E=8$ GPa
4b Reservoir Young’s modulus $E=20$ GPa
5.3.3 5a UGS uneven $\Delta P$: $\Delta P_{1}=-10$ MPa, $\Delta P_{2}=0$ MPa
5b UGS uneven $\Delta P$: $\Delta P_{1}=-10$ MPa, $\Delta P_{2}=-20$ MPa
5.3.4 6a Principal stress $\sigma_{H}$ and $\sigma_{h}$ rotated by $\theta=90^{\circ}$
6b Ratios $M_{1}=\sigma_{h}/\sigma_{v}=0.40$ and $M_{2}=\sigma_{H}/\sigma_{v}=0.47$
Table 2: Scenarios addressed in Stage 1 of the sensitivity analysis.
The main interest of this study is focused on the fault behavior, and in particular the stress conditions yielding a potential sliding. According to the classical Coulomb failure criterion, fault stability is guaranteed if the modulus $\tau$ of the shear stress is smaller than its limiting value $\tau_{L}$:
$\tau<\tau_{L}=c+\mu(\sigma_{n}-P_{f}),$ (1)
with $\mu$ the friction coefficient, i.e., the tangent of the friction angle, $\sigma_{n}$ the normal compressive stress, and $P_{f}$ the fluid pressure in the fault. Wherever $\tau=\tau_{L}$, sliding is allowed with a potential energy release, while the condition $\tau>\tau_{L}$ is forbidden. Three parameters are used to evaluate quantitatively the fault state: (i) the criticality index $\chi=\tau/\tau_{L}$, and (ii)-(iii) the area $t_{a}$ and the sliding $d$ where $\tau=\tau_{L}$. Clearly, from (1) we have $\chi$in [0, 1]. The closer $\chi$ to 1, the more likely is a fault activation. In particular, in the sensitivity scenarios the maximum criticality factor ($\chi_{\max}$), the maximum sliding ($d_{\max}$), and the fault area ($t_{80}$) where $\chi$ exceeds 0.8 are used to lump the faults behavior.
5 Modeling Results
5.1 Fluid-dynamic model
The gas flow dynamic during injection and withdrawal is simulated by different codes according to the selected fluid: OPM Flow (The Open Porous Media Initiative, 2023; Rasmussen et al., 2021) for CH4 (UGS), and Eclipse (Schlumberger, 2014) for H2 (UHS), CO2 (CCS), and N2. For all simulations, the same well configuration is used. During the PP stage, two wells located next to the reservoir boundary are open in the first layer. Then, these wells are converted to injection purposes. Only for UHS, nine storage wells serve for injection and withdrawal. These wells are evenly distributed throughout the reservoir. In Figure 5, the location of the injection and production wells is shown with respect to the fault system. To avoid interpolation between different computational grids, both OPM and Eclipse finite volume cells coincide with the hexahedral elements of the geomechanical model. Each reservoir block is subdivided into 4000 regular cells with a $20\times 20\times 10$ partitioning. The characteristic horizontal and vertical permeability is $k_{h}=600$ mD and $k_{v}=300$ mD, respectively.
Figure 5: Areal location of the injection/production wells (blue and red for injection/production for CH4, CO2, N2 and PP for H2; green for UHS cycles) and example of pressure distribution from a dynamic simulation for CH4 (on top) and H2 on bottom within a reservoir block. The grid is split to provide evidence of the pressure distribution along the vertical direction.
Figure 6: Maximum pore pressure variation in time for the investigated fluids. The leftmost panel is the common pressure decline during the primary production stage.
The injection-production history is depicted in Figure 6. The PP stage is conventionally set to 10 years with a maximum pore pressure drop of about 18 MPa. The targeted final Recovery Factor, i.e., the ratio between the cumulative produced volume of gas and the gas originally in place, is 90%, in line with theoretical reference behavior for a depletion-driven gas reservoir. Then, for CH4 the reservoir pressure recovers to the initial value in two years during the CGI phase and a set of UGS cycles, consisting of 6-month production and 6-month storage each with a 10-MPa pressure excursion, follows. The simulation results show that the pressure change throughout the production (or injection) phase in a single UGS cycle remains nearly uniform in space. Hence, employing a uniform pressure change within each block at every loading step could be also fully warranted. As to the other fluids, the pressure history is targeted to closely approximate the maximum drop in reservoir static pressure observed for CH4, with some adjustments needed to achieve more realistic storage strategies. In particular, the 20-MPa pressure recovery increment during CGI is carried out in a longer period and no production/injection cycles are simulated for CO2 and N2 storage. The CO2 storage scenario considers a maximum injection rate of 4,106 Sm3/day per well for 13 years before recovering the initial pressure value. Injection of N2 lasts 6 years with an injection rate of 4,106 Sm3/day per well. Concerning H2, 10 years of injection/production activity are simulated after a 13-year refilling phase at a maximum rate of 4,106 Sm3/day per well. Each UHS cycle consists of a 6-month withdrawal and a 6-month injection period with a maximum pressure excursion of 9.2 MPa. In all simulations, a 1-year loading step (l.s.) is used during PP and CGI/refilling, while during UGS and UHS the loading step is equal to 15 days.
It is important to recall that in the Netherlands the fluid pressure is not allowed to exceed the natural initial value $P_{i}$. Pore pressure propagation is also accounted for by simulating its variation within a permeable finite-thickness fault zone. The fluid pressure in the fault, $P_{f}$, generates forces that opposes to the action of the pressure variation on $\sigma_{n}$.
5.2 Mechanisms of fault re-activation: reference case (Scenario #1)
The basic fault reactivation mechanisms are studied in scenario #1, which serves as the reference scenario for UGS activity with CH4. A few preliminary outcomes can already be found in Teatini et al. (2019, 2020) and Franceschini et al. (2024). For the sake of completeness, they are briefly summarized here as well.
Figure 7 shows the behavior of the maximum $\chi$ value, $\chi_{max}$, experienced on faults F1 through F5 for 13 years, i.e., PP, CGI and the first UGS cycle. Faults F1 and F2 can reach a critical condition at the end of the PP stage. During CGI and UGS, the fault activation risk reduces, with $\chi_{max}$ decreasing to about 0.80. During CGI, $\chi_{max}$ initially decreases, i.e., the fault is more stable, but then increases again, with the shear stress acting in the opposite direction with respect to that experienced during PP. The time behavior of the maximum sliding $\delta_{max}$ computed for each fault is also shown in Figure 7. These two quantities, i.e., $\chi_{max}$ and $\delta_{max}$, are closely related each other, since a single element can slide only when $\chi=1$. However, we prefer to propose an averaged version of $\chi$, so as to obtain information on the criticality state of the entire fracture at a given depth. In particular, for the sake of clarity and ease of readability, $\chi$ is represented for each fault as a function of depth only, i.e., for each $z$-value we compute the $\chi$ average for the stripe of fault elements located at the same depth. For this reason, it is possible to detect sliding conditions even when the $\chi$ value reported on the top frame of Figure 7 is smaller than 1. Fault F1, F2, F4, and F5 can start sliding between year 4 and 6. The maximum sliding can be of about 1.4 cm. Fault F3 remains unaffected throughout the loading and injection steps due to the model symmetry. The most critical condition develops along the top and bottom of the reservoir. Figure 8 shows the behavior of $\chi$ with respect to depth on all the fracture surfaces at loading step 10.
Figure 7: Scenario #1: $\chi_{max}$ (top), $\delta_{max}$ (middle), and $\tau$ in time for all faults. Positive values for the shear stress means that it is upward oriented.
Figure 8: Scenario #1: $\chi$ in depth at the PP end taken at the red point (on the left sketch).
The mechanisms leading to the possible fault reactivation can be better understood in terms of shear stress $\tau$ (bottom frame of Figure 7), particularly at at loading steps 0, 10 (end of PP), 11, 12 (end of CGI), 12.5 (end of UGS extraction), and 13 (end of UGS injection). The largest values are computed at the end of primary production, with opposing shear stress orientations at top and bottom of the reservoir. During PP, the direction of the shear stress is always oriented toward the reservoir mid-plane. At loading step 4 and 5 sliding starts (middle of Figure 7) at reservoir top and bottom. The elements surrounding the activated fault portions increase their shear stress to accommodate the excess to the Coulomb limit not supported by the activated (sliding) elements. At loading step 11, half of the pore pressure change has been recovered. As the reservoir expands due to pressure recovery, the shear stress changes direction and annihilates approximately the shear stress acting on the previously sliding elements. The reservoir continues to recover pressure and re-expand until loading step 12. During the second half of cushion gas injection the shear stress increases again, with a direction opposite to that experienced during primary production. The expansion during CGI can generate again a critical condition on the fault, with a possible localized reactivation at the reservoir top due the stress redistribution following the sliding. During UGS, at the end of the production phase (loading step 12.5), the shear stress almost equals the condition at loading step 11. Correspondingly, the shear stress at loading step 13 equals the stress state at loading step 12. Indeed, during UGS, sliding does not occur and the porous medium behaves according to a linear elastic constitutive law. This complex behavior is also confirmed by the analysis of the stress path experienced by some representative elements, as reported by Franceschini et al. (2024).
5.3 Sensitivity analysis: Stage 1
Starting from the reference configuration of scenario #1, we explore how variations in individual geometric or mechanical parameters can affect the outcomes. We conducted an extensive investigation, of which we report here the most significant results referring the scenarios summarized in Table 2 for the UGS analysis with CH4. Other simulations were carried out, but are not discussed here for the sake of brevity. The outcomes are cross compared to identify the settings that make the subsurface system more prone to fault reactivation. For the analysis, we use the parameter $t_{80}$, i.e., the areal extent where $\chi\geq 0.80$. Such areal extent is scaled by 2,000 m, which is the characteristic size of each reservoir block, so that $t_{80}$ has the size of a length. For the sake of simplicity, we focus on faults F2 and F3, unless differently specified.
5.3.1 Reservoir and fault geometry
The impact of fault geometry is investigated by varying the dip angle of the central fault F3 (#2a - #2b) and increasing the offset between reservoir compartments of half (#2c) and entire (#2d) reservoir thickness. Figure 9 shows $\chi_{max}$ in time for all faults in all scenarios. Interestingly, while the geometry variation has no effect on fault F1, a large offset between the reservoir compartments (scenario #2d) enhances the potential for instability in faults F4 and F5. The instability of fault F2 is slightly increased by a variation of the fault geometry, however this becomes significant only when a large offset between compartments is introduced (#2d).
Figure 9: Scenarios #2: behavior of $\chi_{max}$ from all the loading steps for each fault. Note that F4 and F5 exhibit an identical behavior, due to symmetric conditions.
Figure 10: Scenario #2d: maximum sliding $\delta_{max}$.
In the latter configuration, unlike the reference scenario, fault F3 is significantly affected by the asymmetry of the system. Indeed, it becomes active and experiences the highest stress levels at the top and bottom of each compartment. Unlike in the reference scenario, now $\chi_{max}$ on fault F3 can reach the limiting value. Furthermore, $\chi_{max}$ achieves the alert threshold of 0.8 also during CGI and UGS, suggesting that UGS operations in this condition might be critical. Consistently, fault F3 exhibits the highest sliding value $\delta_{max}$ of 7.4 cm at the end of the PP stage (Figure 10). Some instability can also occur at the end of the UGS cycle. The other faults can activate around the loading step 4 again at the reservoir top first, and at the reservoir bottom later. Figure 11 shows the active elements within the embedded fault discretization at loading step 6, 10 (end of PP) and 12.5 (middle of a UGS cycle).
Figure 11: Scenario #2d: active interface elements (IE) on the fault system at loading step 6 (left), loading step 10 (center), and loading step 12.5 (right). Only the elements at the reservoir top, and then bottom, are activated.
5.3.2 Geo-mechanical parameters of faults and reservoir
The investigation continues further into the effect of Coulomb rupture criterion and reservoir stiffness. Concerning the Coulomb parameters, the variation of both the cohesion (#3a) and friction angle (#3b) exerts a significant influence on fault stability, mainly during the PP and CGI phases. A reduced friction angle, in particular, weakens the resistance to fault reactivation.
The mechanism of fault weakening is investigated by using slip-weakening constitutive law for the fracture behavior (Franceschini et al., 2024):
$\varphi=\begin{cases}\varphi_{s}+\dfrac{\varphi_{d}-\varphi_{s}}{d_{c}}\delta&\text{for }\delta<d_{c}\\ \varphi_{d}&\text{for }\delta>d_{c}\end{cases}$ (2)
The two parameters defining the new constitutive law are $\varphi_{d}$ and $d_{c}$, i.e., the dynamic friction angle and the slip weakening distance, respectively. Slip weakening occurs after sliding begins by reducing the value of the friction angle. The most critical results in terms of fault stability are obtained for scenario #3c, characterized by a reduction of the friction angle from $\varphi_{s}=30^{\circ}$ to $\varphi_{d}=10^{\circ}$ in a slip distance of $d_{c}=2$ mm. The behavior of $\chi_{max}$ for scenarios #3 during the CGI and UGS stages is shown in Figure 12, showing that slip-weakening may cause a reactivation also at loading steps 11 and 13, but not at loading step 12.5, i.e., in the middle of a UGS cycle.
Figure 12: Scenarios #3: $\chi_{max}$ in time over the CGI and UGS phases.
Figure 13: Scenarios #4: $\chi_{max}$ in time over the CGI and UGS phases.
The relationship between reservoir stiffness and fault behavior has been further investigated in Scenarios #4. A significant contrast between the reservoir and the surrounding caprock, sideburden, and underburden stiffness may play an important role to induce a critical stress state on faults. In scenario #6a the reservoir is softer than the surroundings, and this generally increases the likelihood of a fault reactivation (see Figure 13). Conversely, in scenario #4b the fault reactivation likelihood is notably reduced. During the UGS stage, $\chi_{max}$, remains below 0.7, regardless of the reservoir stiffness value (see Figure 13).
5.3.3 Uneven pore pressure excursion in the reservoir compartments
The effect of a different pore pressure change in the two reservoir compartments is investigated in Scenarios #5 (Table 2). During UGS, fault F3 and, secondarily, faults F4 and F5, may exhibit larger $\chi$ values at loading step 12.5 (Figure 14). When $\Delta P_{2}=0$ MPa, faults F4 and F5 keeps the same critical degree experienced at the end of CGI over the UGS phase. Conversely, with $\Delta P_{2}=20$ MPa the most critical condition is achieved in the middle of the UGS cycle, where $\chi_{max}$ is larger than the threshold value 0.80. Figure 15 shows the distribution of the potentially active elements ($\chi\geq 0.80$) in Scenarios #5, again along the reservoir boundaries for block 2.
Figure 14: Scenarios #5: $\chi_{max}$ in time over the CGI and UGS phases.
Figure 15: Active interface elements (IE) on the fault system at loading step 12.5 for scenario #1 (left), #5a (center) and #5b (right).
For scenario #5b, Figure 16 shows the time behavior of $\chi_{max}$ and the shear stress modulus $\tau$ for a node located at the reservoir bottom on fault F4. The shear stress increases during the PP stage, reaching the limit shear stress $\tau_{L}$ at loading step 5. Then, $\tau$ equates $\tau_{L}$ at the end of CGI and at the end of UGS production and injection cycle. On fault F2 $\chi_{max}$ approaches 1 at the end of the UGS injection phase too, but remains well below the criticality value at the end of CGI phase.
Figure 16: Scenario #5b: $\chi_{max}$ in time on fault F1 (left), $\|\boldsymbol{\tau}\|_{2}$ and $\tau_{L}$ in time on fault F4 (right).
5.3.4 Initial stress regime
The orientation on the principal components of the stress tensor in undisturbed conditions is another key factor for predicting the fault stability during the reservoir management. This is analyzed in Scenarios #6. The modeling outcome revealed that a rotation of the maximum horizontal stress by 90 (scenario #6a) does not provide a significant variation with respect to the reference case. By distinction, an initial stress regime close to normally consolidated conditions (scenario #6b) significantly increases the fault reactivation likelihood, as it can be appreciated from the $t_{80}$ values in Figure 17. Although the occurrence of such a scenario is unlikely, it is important to consider its implications. With the undisturbed stress regime of scenario #5b, the limit $\chi_{max}=1$ is reached at an earlier point of the PP stage on faults F1 and F2 (l.s. 6) and F4 and F5 (l.s. 8), resulting in a larger fault area that is prone to a potential reactivation (Figure 17). The earlier activation during PP causes a reactivation of faults F1 and F2 at the end of the CGI and UGS stages as well.
Figure 17: Scenarios #6: $t_{80}$ in time on fault F2. A zoom for the UGS stage is provided on the right panel of the figure.
5.4 Ranking factors favoring fault reactivation
The objective of this section is to define criteria to rank the potential of inducing “unexpected” seismic events for the different investigated scenarios. The analysis focuses on fault F3, which is the most stressed in the realistic condition of a compartment offset, and fault F2, representing the behavior at the reservoir boundaries. The results are presented in Table 3. The ranking process takes into account the following quantities from the most to the least important one:
  1. 1.
    $\chi_{max}$ during UGS stage;
  2. 2.
    the maximum value of average sliding $\delta_{avg}$;
  3. 3.
    the loading step of first activation.
Fault F2 Fault F3
Scenario $\chi_{max}$ Max Act. $\chi_{max}$ Scenario $\chi_{max}$ Max Act. $\chi_{max}$
UGS $\delta_{avg}$ [m] year UGS $\delta_{avg}$ [m] year
3c 1.00 0.026 7 1.00 2d 0.84 0.033 7 1.00
5b 0.97 0.010 9 1.00 2c 0.81 0.007 - 0.85
6b 0.96 0.018 6 1.00 5a 0.53 $0$ 0.53
2d 0.84 0.008 9 1.00 5b 0.53 $0$ 0.53
3b 0.81 0.010 7 1.00 2a/b 0.32 $0$ 0.46
2c 0.80 0.007 9 1.00 1 0.00 $0$ 0.00
4a 0.79 0.010 8 1.00 6a 0.00 $0$ 0.00
3a 0.78 0.008 8 1.00 6b 0.00 $0$ 0.00
1 0.78 0.007 8 1.00 3a 0.00 $0$ 0.00
5a 0.78 0.007 9 1.00 3b 0.00 $0$ 0.00
2a/b 0.77 0.007 8 1.00 3c 0.00 $0$ 0.00
6a 0.74 0.007 10 1.00 3d 0.00 $0$ 0.00
3d 0.70 0.009 8 1.00 4a 0.00 $0$ 0.00
4b 0.78 0.005 - 0.78 4b 0.00 $0$ 0.00
Table 3: Ranking of the simulated scenarios according to the largest $\chi_{max}$ in the UGS stage, the largest sliding, and the earliest first activation step.
As expected, fault activation during CGI and UGS cycles exhibits distinct critical factors for boundary and central faults. For fault F2, stability is primarily influenced by the initial stress regime of the system, geomechanical properties, like reservoir stiffness, and fault characteristics, such as cohesion, static friction angle, and the presence of slip weakening. In contrast, for fault F3 geometric parameters play a major role. The stability of fault F3 is strongly influenced by a compartment offset, non-vertical fault planes, and differential pressure changes in the two compartments.
5.5 Sensitivity analysis: Stage 2
In the second stage of the sensitivity analysis, we combine the most influential factors for yielding a potential fault reactivation, as identified in Stage 1, and test the new scenarios for the storage of different gases, namely CH4, CO2, H2, and N2. We remember that the primary objective of this analysis is to explain realistic configurations likely to be encountered in the Rotliegend formation, rather than to explore “extreme” conditions. Therefore, taking into account the most influential factors from Stage 1, we chose to focus on the combinations presented in Table 4. They combine a compartment offset, non-vertical fault planes, and variations in the fault characteristics, i.e., cohesion (C1) and static friction angle (C2). Two dedicated combinations analyze the effect of different geomechanical properties, i.e., decreasing (W) and increasing (H) of the reservoir stiffness.
Combination Parameter/mechanism
C1 $c=0$ MPa, $\delta=+65^{\circ}$, $o=100$ m
C2 $\varphi_{s}=20^{\circ}$, $\delta=+65^{\circ}$, $o=100$ m
W $E$ decreased by 30% during injection
H $E$ increased by 30% during injection
Table 4: Combination of settings addressed in Stage 2 of the sensitivity analysis for the fluids under consideration.
The potential risk of fault reactivation is evaluated for each scenario by the parameter $t_{80}$, whose behavior in time is shown in Figure 18. Regardless of the fluid, Figure 18 shows that the combination C2 provides the realistic conditions yielding the larger $t_{80}$ values. This scenario is characterized by a compartment offset and a reduced friction angle. Conditions close to reactivation occur during both the CGI and the UGS/UHS, or long-term storage, stages, especially if the faults already experienced a reactivation during the PP. As compared to other critical combined configurations, such as C1, scenario C2 is more critical primarily because of the reduction of the friction angle.
Figure 18: Sensitivity analysis, Stage 2: $t_{80}$ in time for fault F2 and F3 and the different fluids under investigation.
Because of its significance, the outcomes of the critical combination C2 have been separately described for CH4 and the other gases. This distinction is made due to the differences in pressure history (see Figure 6).
5.5.1 UGS critical scenario
The combination of factors in configuration C2 causes a potential reactivation mainly during primary production. The condition is particularly critical for faults F2 and F3 during the UGS phase, with both faults reaching a critical conditions not only at the end of the UGS storage phase (UGS-S) but also during withdrawal (UGS-P). Figure 19, which shows $\chi_{max}$ in time for scenario C2 as compared to the reference case (scenario #1), highlights that critical conditions for fault reactivation develop earlier during PP. $\chi_{max}$ is also greater than, or close to, 0.8 during CGI and UGS. Because of the loss of symmetry, F2 is more stressed than F1 and $\chi_{max}$ achieves values larger than 0.8 on fault F3 as well. Figure 20 provides a 3D view of $\chi$ distribution in space at a few significant loading steps.
Figure 19: Scenario C2: $\chi_{max}$ in time for all faults as compared to the reference case (scenario #1).
Figure 20: Scenario C2: $\chi$ distribution in space on the fault system at the end of PP (left), CGI (middle) and UGS (right).
The behavior of the maximum displacement ($\delta_{max}$) for all the faults is shown in Figure 21. During the CGI phase, all the faults slide, primarily during the period between loading step 11 and 12, in the opposite direction to what occurred during primary production. Additionally, faults F1, F2, and F3 slightly slide again at the end of the UGS injection phase.
Figure 21: Scenario C2: maximum sliding $\delta_{max}$ in time.
5.5.2 CO2, H2, N2 storage dynamics and fault activation
The goal of this section is to compare how the temporary or permanent storage of different gases may influence the fault stability as measured by $\chi_{max}$ in the critical scenario C2. Since the pore pressure variation during PP is the same for each fluid, $\chi$ behaves identically until l.s. 10. After that, the impact on fault stability must be considered based on whether temporary or permanent storage scenarios occur.
For the permanent storage scenarios, i.e., CO2 and N2, at the beginning of the storage $\chi_{max}$ decreases due to the unloading of the shear stress caused by reservoir expansion. However, as the storage continues, the shear stress is loaded with the opposite sign, reaching a relative maximum at the end of the injection stage. This point corresponds to the full pore pressure recovery (Figure 22). As far as it concerns the storage/production cycles, $\chi_{max}$ on F1, F2 and F4-F5 reaches its maximum at the end of the UHS injection stage, corresponding to the highest value of pressure in the reservoir. For F3, $\chi_{max}$ exhibits an additional peak at the end of the UHS withdrawal stage.
Figure 22: Scenario C2: $\chi_{max}$ in time for CO2, N2 and H2.
5.5.3 Geo-chemical effects on fault stability
The inclusion geochemical effects is also considered. Based on the literature review summarized in Section 3, we have incorporated the possibility to account for geochemical influences by introducing a hardening/weakening behavior of reservoir rocks. Specifically, during the injection the Young modulus is increased or decreased by 30%, depending on the fluid involved. For CO2, both weakening and hardening scenarios have been observed, depending on the actual rock mineralogical composition. Conversely, H2 induces a weakening effect, while N2 shows a relatively neutral geo-chemical impact on the reservoir rock.
Figure 23 shows the behavior of $\chi_{max}$ for the weakening/hardening scenarios (W/H). Fault F3 remains stable throughout the simulation because of symmetry. During CO2 injection with a hardening behavior, $\chi_{max}$ first decreases and then increases with respect to the reference case. The opposite occurs for the weakening scenario. Similar considerations hold for H2 in the weakening scenario, with an even smaller difference with respect to the reference case. Based on these results, we can preliminarily conclude that geochemical effects on the reservoir rock have a minor impact on fault stability.
Figure 23: Scenario W/H: $\chi_{max}$ in time for CO2 (upper) and H2 (lower).
6 Discussion
A few papers addressing the topics of this project have recently been published. The papers by Wassing et al. (2017) and Haug et al. (2018) are of particular interest because they focus on Rotliegend reservoirs in the Netherlands and northern Germany, providing insights into fault reactivation and fault rupture during primary production. Although the authors: i) do not consider induced seismicity during UGS stages, ii) use a simplified 2D setting, and iii) employ a different modeling approach concerning fault activation, these studies carry out a parametric analysis on the same geometric/geomechanical features investigated in this project. Therefore, it is interesting to see if our results are consistent with those presented in these publications.
The findings of our work mostly align with the observations by Wassing et al. (2017) regarding fault failure mechanisms and slip initiation. Our study confirms that fault slip initiates at the top of the reservoir rocks where the initial shear capacity utilization SCU, corresponding to our safety factor $\chi$, is maximum. Both studies agree that faults without offset experience reactivation at later stages of depletion, and fault rupture does not extend upwards into the caprock. Scenarios #2c, #2d, and C2, for example, confirm that in the case of a block offset, an earlier fault activation is favored and remains bounded within the reservoir depth range. Regarding the influence of contrast in elastic properties distribution, our study partially agrees with Wassing et al. (2017). They observed a smaller peak in shear stress at the bottom of the footwall reservoir block and a localized decrease in shear stresses and SCU at the bottom of the hanging wall reservoir block. From a qualitative point of view, we agree. It must be considered that Wassing et al. (2017) assigned a uniform density and elastic properties to all rocks. Our study, in particular scenario #4a and #4b, demonstrates that the contrast in elastic properties may significantly impact on stress concentration and, therefore, fault failure.
Comparing our results with the findings presented by Haug et al. (2018) regarding factors influencing fault reactivation and criticality, we agree that the initial stress regime (scenario #6a and #6b) play significant roles in fault behavior. Indeed, fault criticality largely increases as the horizontal components of the natural stress regime decrease relative to the vertical stress. Furthermore, we confirm that the stiffness contrast between the reservoir and surrounding rocks (scenario #4a and #4b) governs the distribution of production-induced loading and the degree of stress rotation, impacting fault reactivation with pressure depletion. We demonstrated that stiffness contrast between rocks is a significant factor in stress redistribution, favoring fault reactivation when the contrast increases. According to Haug et al. (2018) the depletion of thicker reservoir horizons results in a stronger fault-loading compared to the depletion of thinner reservoir horizons, as the thick reservoir undergoes a relatively larger strain for the same pore fluid pressure decrease. Although our sensitivity analysis did not directly consider reservoir thickness as a parameter, we observed consistent outcomes when the larger strain is attributed to a larger pressure decrease (scenario #5b).
However, there are a few areas where our project disagrees with findings from Haug et al. (2018). While they proposed that the fault throw should be half of the reservoir thickness to obtain maximum shear stress ratio (SSR) values, we found that the most critical condition occurred when the fault offset is equal to the entire reservoir thickness. These discrepancies can be attributed to variations in the modeling setup, including the horizontal-to-vertical ratio of the natural stress components and the reservoir depth. These disparities emphasize the importance of modeling approaches and setup parameters and warrant further researches to fully understand faults behavior in different geological contexts.
6.1 Definition of safe operational bandwidths
After the end of primary production, when the natural fluid pressure $P_{i}$ is reduced to $P_{min,PP}$, the reservoir experiences a relatively fast pressure recovery to $P_{max,CGI/ST}$ during cushion gas injection or gas storage (CO2 or N2). Afterward, UGS or UHS are characterized by a cyclic pressure fluctuation between $P_{min,UGS/UHG}$ and $P_{max,UGS/UHS}$ at the end of the production and injection phases, respectively. In the usual practice and in alignment with the legislation of some countries, e.g., the Netherlands, $P_{max,CGI/ST}\simeq P_{max,UGS/UHS}\simeq P_{i}$ and $P_{min,UGS/UHS}\geq P_{min,PP}$ (Figure 24a). In this framework, guidelines for the definition of “safe operation bandwidths” in gas storage, i.e., operations with a reduced risk of fault reactivation, must identify proper values for the aforementioned pressure bounds. Nevertheless, it has to be recalled that a fault reactivation could always occur aseismically (Cappa et al., 2019) .
The interpretation of the modeling results allows to outline some key guidelines. They are necessarily qualitative because of the theoretical/general framework of the modeling application and the quasi-static nature of the implemented model, which properly simulates the possible inception of fault slip but not seismicity. The numerous scenarios investigated within the study have clearly revealed that fault failure is more likely to happen during CGI/ST and UGS/UHS in depleted reservoirs when:
  1. 1.
    fault reactivation is occurred during PP at pressure $P_{seis,PP}$ (Figure 24);
  2. 2.
    the reservoir pressure approaches $P_{max,CGI/ST}$, $P_{max,UGS/UHS}$, or $P_{min,UGS/UHS}$.
The resulting guidelines, which link $P_{max,CGI/ST}$, $P_{max,UGS/UHS}$, and $P_{min,UGS/UHS}$ to $P_{min,PP}$ and $P_{seis,PP}$, can be summarized as follows.
Figure 24: (a) Pressure values to be accounted for in the definition of operational safety guidelines. (b) and (c) Safe pressure variation during CGI/ST and UGS/UHS phases in the case of a seismic event occurring during PP at pressure close to $P_{i}$ and $P_{min,PP}$, respectively.
  • 1.
    If a fault reactivation occurs during PP, the pressure change $\Delta P$ spanned during CGI/ST and UGS/UHS phases should be kept smaller than $|P_{i}-P_{min,PP}|$. Indeed, a number of investigated scenarios (see for example Figure 9) reveals that fault activation during primary production leads to a stress redistribution and a new deformed “equilibrated” configuration that is newly loaded, in the opposite direction, when the pressure variation changes the sign. A reasonable rule is to keep $\Delta P$ smaller than maximum between $|P_{seis,PP}-P_{min,PP}|$ (case a) and $|P_{i}-P_{seis,PP}|$ (case b). Generally, case a) or case b) prevails depending on the value of $P_{seis,PP}$ with respect to $P_{i}$ and $P_{min,PP}$. (Figure 24b,c).
  • 2.
    In case a, $P_{max,CGI/ST}$ and $P_{max,UGS/UHS}$ should be kept below $P_{i}$. A reasonable rule of thumb could be to keep $P_{max,CGI/ST}$ and $P_{max,UGS/UHS}$ smaller than the $P_{seis,PP}$.
  • 3.
    In case b, $P_{min,UGS/UHS}$ should be kept above $P_{min,PP}$. A reasonable rule of thumb could be to keep $P_{min,UGS/UHS}$ larger than the $P_{seis,PP}$.
  • 4.
    If during PP activation occurs on a fault separating two reservoir compartments, during CGI/ST or UGS/UHS the pressure difference between the adjacent blocks should be safely kept smaller than $P_{i}-P_{seis,PP}$.
  • 5.
    If no activation occurs during PP, $P_{max,CGI/ST}$ can equate $P_{i}$ with no particular risk of unexpected events during CGI and UGS. Moreover, $\Delta P$ during UGS/UHS can safely span the whole pressure change between $P_{i}$ and $P_{min,PP}$. In fact, the system lies in reloading conditions and behaves practically in an elastic way within the pressure range already experienced.
It has to be emphasized that fault reactivation is largely case-sensitive, depending on the geometry of the fault/reservoir (e.g., the presence of sloped faults, dislocation of the reservoir compartments), differential pore pressure between adjacent reservoir compartments and within each reservoir block, the geomechanical properties of the faults and the reservoir, caprock, under- and overburden. Therefore, specific investigations are fundamental to characterize the actual reservoir setting and quantify more specific bounds. Moreover, the need of more specific analyses holds in relation to the geochemical effects on fault mechanical properties caused by the substitution of formation fluids with CO2, H2 or N2. In the modeling approach used in this work, these possible effects on faults have been neglected due to the lack of laboratory testing. Finally, we underline that the above rules are aimed at avoiding, or at least limiting the probability, of fault activation. For example, according to Wassing et al. (2017), “after the onset of fault reactivation, a further pore pressure decline of 3.7 MPa (or 1.6 MPa with no fault offset) and a critically stressed length of approximately 30 m is needed for the nucleation of a seismic event”. Therefore, since fault reactivation can develop with aseismic slip, these can be regarded as conservative recommendations.
7 Conclusions
Depleted gas reservoirs serve not only as storage facilities for natural gas but also for CO2, N2, and H2. The injection and extraction of various fluids induce changes in the hydraulic and mechanical state of the reservoir at depth. Consequently, deformations of various magnitude occur beneath the subsurface, potentially leading to induced seismic events along fault planes. While the majority of human-induced seismic activities can be explained by fluid injection at pressures beyond the initial levels, triggering the shear stress along fracture surfaces to reach the limit strength, certain recorded events do not fall into this explanation. These particular seismic events, usually called “unexpected”, occur when the pressure falls within the range experienced during primary production.
The main aim of this work is to understand whether and under which circumstances faults can be reactivated (and seismic events possibly induced) during UGS activities. A one-way coupled strategy is adopted to deal with the fluid-structure interaction. The study has been carried out using the typical geological setting of UGS reservoirs in the Netherlands and four underground storage systems, i.e., cyclic storage of CH4 and H2 and permanent storage of CO2 and N2, have been analyzed. This objective is achieved by investigating which are the main factors that control the reactivation of faults and seismicity under permanent or cyclic storage conditions of the stored gases, and to which extent the mechanisms for seismicity are similar to the mechanisms during natural gas storage. Additionally, the study evaluates the potential risk posed by geochemical effects.
The effects of CO2 injection on mechanical properties of faults are negligible as both weakening and hardening effects have been simulated on reservoir and caprock with a ±$\pm 30\%$ change in stiffness moduli with no significant change in fault stability. Instead, the fault critical conditions are controlled by the extreme values of pore-pressure (minimum and maximum) in the formation. Since these values are similar for the natural gas storage operations, the mechanisms for reactivation (which can be aseismic) are also similar. The critical factors (drivers/settings) influencing fault activation during CGI and UGS cycles arrange differently for the boundary and central faults. On faults bounding the reservoir, during UGS the stability is mainly jeopardized by an initial stress regime with small horizontal principal components, failure criterion with low friction angle, large pressure change because of production/injection, and significant contrast between the reservoir and the over- side- and underburden stiffness. If the central fault is on vertical position and allows for symmetry conditions, the major influencing drivers to fault instability are given by geometrical parameters characterizing the fault/reservoir system. Offset of the reservoir compartments, non-vertical fault plane, different pressure change in the two compartments, and fault dip are the features threatening the stability of this fault. Finally, it is worth to highlight the lack of quantitative data regarding N2 and H2 effects on rock properties, and thus on fault stability.
These considerations are supported by the ranking of the combined scenarios. A re-assessment of the project’s findings on geochemical effects causing mechanical weakening should be made at the point where more data from laboratory experiments and case studies becomes available.
CRediT authorship contribution statement
Selena Baldan: Investigation, Formal analysis, Writing - original draft. Massimiliano Ferronato: Conceptualization, Methodology, Writing - review and editing, Supervision. Andrea Franceschini: Methodology, Software, Writing - review and editing. Carlo Janna: Software. Claudia Zoccarato: Conceptualization, Methodology, Formal analysis, Investigation. Matteo Frigo: Software, Investigation. Giovanni Isotton: Conceptualization, Methodology, Formal analysis, Software, Funding acquisition. Cristiano Collettini: Formal analysis, Writing - review and editing. Chiara Deangeli: Methodology, Formal analysis, Writing - review and editing. Vera Rocca: Methodology, Formal analysis, Writing - review and editing. Francesca Verga: Methodology, Formal analysis, Writing - review and editing. Pietro Teatini: Conceptualization, Methodology, Formal analysis, Writing - review and editing, Supervision, Funding acquisition.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This research was supported by the State Supervision of Mines (SodM), Ministry of Economic Affairs (The Netherlands), projects KEM01 “Safe Operational Bandwidth of Gas Storage Reservoirs” and KEM39 “Study within the Mining Effects Knowledge Program (KEM-39) on the cyclic storage of gases in the Netherlands”. M.F., A.F., and C.J. from the University of Padova are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). Computational resources were provided by University of Padova Strategic Research Infrastructure Grant 2017: “CAPRI: Calcolo ad Alte Prestazioni per la Ricerca e l’Innovazione”.